A photographic film is coated with a silver bromide layer. When light falls on this film, silver bromide molecules dissociate and the film records the light there. A minimum of 0.6 eV is needed to dissociate a silver bromide molecule. Find the maximum wavelength of light that can be recorded by the film.

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. w base 0 = 0.6 ev For w0 to be min ‘λ’ becomes maximum. W base 0 = hc/λ or λ = hc/w base 0 = 6.63 * 10^-34 *3 * 10^8/0.6 * 1.6 * 10^-19 = 20.71 * 10^-7 m = 2071 nm